OpenStudy (anonymous):

estimate the solution to the linear system 2x+y=-1 5x-3y=1

OpenStudy (anonymous):

What would you do first?

OpenStudy (anonymous):

you can use substitution or elimination. Which do you prefer?

OpenStudy (anonymous):

elimination

OpenStudy (anonymous):

first pick a variable to eliminate so which variable do you want to eliminate first?

OpenStudy (anonymous):

y

OpenStudy (anonymous):

to eliminate for y you could multiply the first equation by 3 since 3 is the common number for 1y and 3y, then add them so 3(2x+y=-1) 5x-3y=1 6x+3y = -3 5x-3y = 1 (+) ------------------- 11x + 0y = -2 11x = -2-- divide both sides by 11 x = -2/11

OpenStudy (anonymous):

do the same process as ^^ to solve for y first eliminate the x then solve for y

OpenStudy (anonymous):

yeah thanks \[thank you \]

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

\[FAN N MEDAL\]

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

\[:)\]